# ZeroDL: Zero-shot Distribution Learning for Text Clustering via Large Language Models

### v1

@@ 4 Experiments
AutoL(Best) 86.8 84.9 41.8 38.5
AutoL(Worst) 82.5 80.7 41.7 38.6
@@
ZeroDL(Mistral) 82.4 79.4 43.4 36.0
ZeroDL(GPT3.5) 73.3 79.1 42.8 42.1
Table 3: The performance with various class labels. The best AutoL in SST-2 are [Wonderful, Bad] and the worst labels are [Irresistible, Pathetic]. In SST-5, [Terrible, Better, Good, Extraordinary, Unforgettable] are the best while [Awful, Better, Hilarious, Perfect, Wonderful] are the worst. GPT3.5 with ZeroDL generates only 4 classes in SST-5.
We evaluate the model performance using accuracy. To determine the predicted class, we leverage the Class n: anchor tokens within the LLM outputs. LLMs might not directly predict the same classes as the ground-truth labels, for example, Class 0 means Positive in the prediction while Negative is labeled as Class 0 in the ground-truth. We thus test all possible mapping combinations and report the best performed mapping. However, this results in a total of factorial of C potential mappings, so we split datasets with more than 7 classes (i.e., DBpedia and YahooAnswers [^3] After removing the 3 smallest size of classes: CarsAndTransportation, SocialScience, Sports) into 2 subsets Front (F) and Back (B) to avoid out-of-memory.
5 Results
Table 1 reports the performance of our method. Interestingly, ZeroDL achieves even better performance than models provided with ground-truth (i.e., Gold) class labels on several datasets. This suggests that ZeroDL might uncover richer or more nuanced class structures within the data compared with the pre-defined labels. Our method demonstrates comparable performance to K-Means clustering that utilizes LLM embeddings [^4] However, llm2vec involves additional training steps., outperforming on tasks with relatively smaller datasets. ZeroDL achieves this by flexible zero-shot prompting. Table 8 in the Appendix shows the importance of the constraints utilized by ZeroDL with detailed examples.
Table 3 investigates the significance of class labels in text clustering tasks. We explore the performance of mistral-7b-instruct-v0.2 using class labels suggested by AutoL Gao et al. (2021) designed for prompt-based model fine-tuning. These labels represent a curated selection. We also investigate the class labels generated by gpt-3.5-turbo. The results demonstrate that clustering performance is generally higher when using manually selected labels compared with the original dataset labels. This suggests that carefully chosen labels can significantly improve clustering outcomes. In the context, ZeroDL performs particularly well on SST-5, implying that the critical role of class labels in text clustering and the potential of ZeroDL to capitalize on informative class labels automatically.
Table 2 shows the examples of class labels generated by ZeroDL. Notably, these labels provide richer explanations for the classifications compared with the original ground-truth labels [^5] The generation of class description might depend on datasets and LLMs; Llama-3 usually does not make it.. Table 6 in the Appendix delves deeper into the data points classified into newly emerged classes, such as "Ambiguous Sentiment" and "Mixed Sentiment," which ZeroDL might uncover based on the data.
ZeroDL involves a trade-off in computational cost. To investigate this aspect, we measure the performance changes based on the amount of input data used (see Table 5 in the Appendix). The results indicate that using only 10% of the data yields plausible performance but it leads to inconsistency in the model performances, considering increased standard deviation.
@@ 6 Conclusion
We introduce ZeroDL, a novel approach to contextualize LLM tasks for a given LLM. ZeroDL employs open-ended zero-shot inference and output aggregation to learn data distributions. We demonstrate its effectiveness, showing competitive performances against embedding-based clustering methods and superior performance than ground-truth labels in some cases. Beyond its clustering capabilities, ZeroDL offers the generation of informative class labels that provide deeper insights into LLMs.
@@ 7 Limitations
Prompt Dependency and Heuristics.
@@
Experiments with Diverse LLMs and Prompts.
While we acknowledge the computational limitations (and price) of ZeroDL, investigating its behavior with a wider range of LLMs (including commercial models like GPT-4, Claude, and Gemini) and prompt templates could provide valuable insights into the generalizability and robustness of the approach.
Lower than State-of-the-Art Performance.
@@
Although ZeroDL have an alternative approach to reduce computational burden through data sampling, effectively sampling data to generate appropriate class labels remains a challenge. The technique within the framework presents a valuable future direction.
@@ Appendix A Appendix
Yahoo(F) 59,518 8,879 10,489 7 ArtsAndHumanities, BeautyAndStyle, BusinessAndFinance, ComputersAndInternet, ConsumerElectronics, EducationAndReference, EntertainmnentAndMusic
@@
[C-T] IMDB SST-2 SST-5 YRev AGNew DBp(F) DBp(L) Yah(F) Yah(L) Macro Micro
All 85.3 82.4 43.4 51.6 75.4 64.0 73.3 47.2 69.4 65.8 65.1
(std) 3.40 3.67 2.93 1.35 2.23 3.17 3.82 1.02 2.65
1% 75.2 70.7 38.3 51.1 - 48.9 37.9 23.7 17.2 45.4 47.8
(std) 23.01 13.47 6.17 1.30 - 10.51 22.91 1.82 2.50
5% 92.9 82.4 45.8 52.1 - 59.2 58.5 35.2 51.7 59.7 60.1
(std) 0.30 2.34 2.78 1.64 - 6.77 7.37 9.03 14.15
10% 91.9 82.0 35.0 50.3 - 67.2 68.6 37.3 57.1 61.2 63.5
(std) 0.82 3.17 4.45 3.47 - 6.35 3.27 2.47 6.24
[T-C] IMDB SST-2 SST-5 YRev AGNew DBp(F) DBp(L) Yah(F) Yah(L) Macro Micro
All 92.2 79.4 36.0 48.2 81.0 62.2 78.4 51.8 73.4 67.0 66.4
(std) 2.34 5.12 3.84 3.22 1.81 4.65 7.14 1.07 1.91
1% 76.1 78.3 39.4 50.0 - 49.3 34.3 21.0 17.2 45.7 46.9
(std) 23.88 6.01 5.17 2.48 - 13.03 19.56 3.67 2.50
5% 93.3 81.0 41.7 48.4 - 55.3 67.9 56.3 55.7 62.5 61.7
(std) 0.93 2.59 4.09 4.98 - 1.84 4.28 6.09 12.68
10% 92.7 74.2 36.9 49.3 - 63.2 75.8 41.8 64.6 62.3 64.6
(std) 1.77 13.36 6.03 5.98 - 6.53 6.57 4.96 3.44
Table 5: ZeroDL performances according to the number of training data. ZeroDL fails to find exact number of clusters in AGNews dataset.
@@
Mixed Sentiment: (...) VeryNeg It ’s hard not to feel you ’ve just watched a feature-length video game with some really heavy back story .
@@
Ambiguous Sentiment: (...) VeryNeg It ’s difficult to say whether The Tuxedo is more boring or embarrassing – I ’m prepared to call it a draw .
@@
Table 8: The examples of generated class labels when no constraints are given; we experiment with wrong task type and unlimited the number of clusters.

### v2

@@ 5 Results
AutoL(Best) 86.8 84.9 43.9 40.5
AutoL(Worst) 82.5 80.7 41.1 39.7
@@
ZeroDL(Mistral) 80.6 83.7 40.3 40.5
ZeroDL(gpt-4.1) 77.0 82.5 41.0 40.3
Table 3: The mistral-7b performance with various class labels. ZeroDL(Mistral) uses only the generated class titles for fair comparison. The best labels from AutoL in SST-2 are [Wonderful, Bad] and the worst are [Irresistible, Pathetic]. In SST-5, [Terrible, Better, Good, Extraordinary, Unforgettable] are the best while [Awful, Better, Hilarious, Perfect, Wonderful] are the worst. gpt-4.1-mini generates [Positive, Negative] (the same as Gold) and [Positive, Negative, Neutral, Mixed, Other / Unclear], respectively.
@@ 6 Conclusion
We introduce ZeroDL, a novel approach to contextualize tasks for a given LLM. ZeroDL employs open-ended zero-shot inference and output aggregation to learn data distributions. We demonstrate its effectiveness, showing competitive performances against embedding-based clustering methods and superior performance than ground-truth labels in some cases. Beyond its clustering capabilities, ZeroDL offers the generation of informative class labels that provide deeper insights into LLMs.
@@ 7 Limitations
Prompt Dependency and Heuristics
@@
Experiments with Diverse LLMs and Prompts
While we acknowledge the computational limitations (and price) of ZeroDL, investigating its behavior with a wider range of LLMs (including commercial models like GPT, Claude, and Gemini) and prompt templates could provide valuable insights into the generalizability and robustness of the approach.
Lower than State-of-the-Art Performance
@@
Although ZeroDL have an alternative approach to reduce computational burden through data sampling, effectively sampling data to generate appropriate class labels remains a challenge. Although we report the trade-off in computational cost and performance in Table 7 of the Appendix, the technique within the framework presents a valuable future direction.
Acknowledgement
The authors would like to thank reviewers in several rounds of discussion. Thanks to Alice Lee for her help with English writing and Minji Hong for her support in writing this work. Lastly, we extend our gratitude to Hyperscale AI Team at NAVER Cloud, for unwavering support in connecting us with key resources throughout this research.
@@ A.1 Data Information
Yahoo(F) 59,518 8,879 10,489 7 ArtsAndHumanities, BeautyAndStyle, BusinessAndFinance, ComputersAndInternet, ConsumerElectronics, EducationAndReference, EntertainmentAndMusic
@@
A.2 Additional Model Performances
TF-IDF KMeans 52.1 53.0 26.2 35.2 43.8 55.9 62.8 44.7 46.5 46.7 48.8
SBERT KMeans 65.0 54.0 33.9 30.3 82.7 95.4 93.9 67.2 68.3 65.6 67.5
llama-3.1 ZeroDL(C-T) 76.0 73.8 35.0 51.1 54.5 37.7 58.5 46.4 47.9 53.4 53.2
-8b-it IDAS(C-T) 92.2 73.8 42.4 45.4 55.4 40.8 64.1 29.2 35.3 53.2 53.9
Gold(C-T) 80.4 73.5 44.6 48.7 64.1 74.4 80.8 49.6 69.5 65.1 66.7
ZeroDL(T-C) 94.4 79.4 39.1 49.4 77.6 46.1 68.1 53.7 60.0 63.1 61.1
IDAS(T-C) 94.6 84.5 38.4 50.1 67.7 37.6 66.4 32.0 36.2 56.4 56.2
Gold(T-C) 94.4 81.5 43.4 50.1 78.9 80.8 75.1 62.8 78.3 71.7 71.2
Qwen-2.5 ZeroDL(C-T) 94.5 90.2 44.5 62.6 71.3 60.8 64.4 53.4 71.3 68.1 67.5
-72b-it IDAS(C-T) 95.3 88.8 47.1 50.5 68.1 62.3 57.2 35.7 55.8 62.3 61.1
Gold(C-T) 95.1 89.4 51.0 57.7 81.3 94.2 97.9 69.0 84.8 80.0 81.6
ZeroDL(T-C) 95.3 90.1 45.4 63.3 76.6 65.2 65.4 55.1 73.6 70.0 69.4
IDAS(T-C) 95.9 89.9 46.3 51.2 76.4 63.2 58.9 34.5 47.7 62.7 61.7
Gold(T-C) 95.8 90.0 52.7 59.9 83.9 96.8 99.7 73.7 87.4 82.2 83.8
Table 5: The performance of ZeroDL for text clustering compared with more methods. C-T denotes the prompt order with class information then input text. T-C is the reversed. Bold means the best accuracy in the same prompt order and underline the outperforming cases than baseline except for ground-truth class labels.
A.3 Qualitative Examples
@@
Mixed Sentiment VeryNeg It ’s hard not to feel you ’ve just watched a feature-length video game with some really heavy back story .
@@
Ambiguous Sentiment VeryNeg It ’s difficult to say whether The Tuxedo is more boring or embarrassing – I ’m prepared to call it a draw .
@@
A.4 Data Sampling Ablation
[C-T] IMDB SST-2 SST-5 YRev AGNew DBp(F) DBp(L) Yah(F) Yah(L)
All 88.8 85.7 40.5 48.7 75.2 58.9 64.1 47.2 61.4
(std) 3.20 0.18 1.22 0.21 1.22 7.85 6.61 1.02 1.92
1% 81.3 79.3 46.9 51.5 66.0 59.4 58.4 36.1 43.7
(std) 2.97 1.5 0.88 0.12 5.71 1.93 5.11 0.72 3.35
5% 74.6 83.6 45.8 49.4 74.2 58.8 66.4 39.3 58.5
(std) 18.49 1.37 2.78 0.47 1.94 5.38 9.13 4.82 2.73
10% 87.0 84.5 44.8 49.2 74.5 67.2 63.9 40.4 60.4
(std) 5.90 1.50 5.02 0.29 1.44 6.35 8.95 5.11 2.96
[T-C] IMDB SST-2 SST-5 YRev AGNew DBp(F) DBp(L) Yah(F) Yah(L)
All 90.2 84.2 36.0 46.8 79.5 56.7 72.2 51.0 66.3
(std) 3.52 0.24 3.84 0.47 1.73 9.73 6.86 0.75 2.80
1% 88.4 58.6 42.0 49.8 74.6 57.7 67.5 37.7 49.4
(std) 0.51 11.93 3.99 2.05 4.34 1.31 8.37 4.1 3.15
5% 85.5 80.2 41.7 46.2 80.7 57.1 74.1 37.2 68.1
(std) 7.36 1.92 4.09 1.42 2.83 10.58 9.36 6.87 4.58
10% 91.2 82.0 39.1 46.6 79.9 63.2 75.3 45.2 68.5
(std) 2.16 1.95 2.26 0.33 2.74 6.53 9.91 3.43 3.42
Table 7: ZeroDL performances according to the number of input data.
A.5 Additional Examples of Generated Classes
@@
Table 9: The examples of generated class labels when no constraints are given; we experiment with different task type and unlimited the number of clusters.